For the following functions, determine minimal SOP realizations:
a. F(a, b, c) = ∑ (0, 1, 2, 3, 4, 5, 6, 7)
b. F(a, b, c) = ∑ (1, 2, 3, 4, 5, 7)
c. F(a, b, c) = ∑ (0, 2, 4, 6)
Answer #1

* Given that f(a,b,c) = E (0, 1, 2, 3, 4, 5, 6,7) ( с таъс + abc Wirabê ( pia mabēt a c + ābēt + abĆ + abc F = ā5 (6+c) +ãb (( Karnaugh Map verification: Method) ! be 10 a of 8 Shells) F=1 (Grosuping b c tabč tabc + abc b) f(a,b,c) : <m (1,2,3,4,5,7)verification: 79 F = 0 + Ⓡ + ® C + 2 b + ab F F = (a b) to {m (0,2,4,5) J Fla, b, c) © 290 +290 +290 +292 - worden toto obsodverification : 11 10 \bc 00 01 . FEĞ bor Exprenions: the following Sop realizatiom Minimum a) f(a,b,c) = {m (0,1,2,3,4,5,6,7)

If you like the answer please thumbs up......i have clarly mentioned all the steps ........